OpenStudy (anonymous):

Perform the multiplication. Assume that no denominators are 0. ((k^2-kg+g^2)/(k^3+g^3))((kc+km+gc+gm)/(c^2-m^2))

OpenStudy (anonymous):

\[\frac{\left(g^2-g k+k^2\right) (c g+c k+g m+k m)}{\left(g^3+k^3\right) \left(c^2-m^2\right)} \]

OpenStudy (anonymous):

\[\frac{1}{c-m} \]Do you want to see the work? The expansion is the sum of four somewhat complex fraction.

OpenStudy (anonymous):

its alright, it would be kinda hard to do it on here

OpenStudy (anonymous):

I will attach a PDF file.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Used Mathematica 8.04 to do the calculations.